OpenStudy (anonymous):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

OpenStudy (raden):

5x + 2y = 18 3x - 7y = -22

OpenStudy (anonymous):

right rad

OpenStudy (raden):

actually, if the choices the given u can use the trials error can u check for the last one, woulde be the solution or not ?

OpenStudy (anonymous):

what would it be?

OpenStudy (raden):

let's check it (if it is D, 2 and 4) for the 1st equation : 5x+2y=18 5(2) + 2(4) = 18 10 + 8 = 18 RIGHT

OpenStudy (raden):

for the 2nd equ. 3x-7y=-22 3(2)-7(4)=-22 6-28=-22 RIGHT TOO :)

OpenStudy (raden):

so, it is D